Hi. I have a question on my homework that I have no idea how to solve. Okay:
Write the equation of each line using the given informaton:
(1/2,3),(1,11/2)
The / sign is not used as a division sign in this problem. Those are fraction. Can you show me how in the world I can solve this?

Answer by stanbon(57400)

Write the equation of each line using the given informaton:
(1/2,3),(1,11/2)
------------------------
Form: y = mx + b
You need to determine m and b.
-------------------------------------
m = (11/2 - 3)/(1 - 1/2) = (5/2)/(1/2) = 5
-----------------------------
You now know m and you know y = 3 when x = 1/2; solve for b.
3 = 5(1/2) + b
---
3 = (5/2) + b
b = 1/2
----------------
Equation:
y = 5x + (1/2)
===============================
Cheers,
Stan H.